# The adjoint state method for parametric definable optimization without smoothness or uniqueness

### 27 Mar 2026 v1

@@ Appendix A Technical results
The following lemma is an immediate consequence of the stronger Aubin property result given in [32, Corollary 4.5]. We give its quick derivation in the proof just for completeness.
@@
The set-valued map of feasible points $R$ is inner semicontinuous, i.e., for all $(x^{*},\theta^{*})\in\mathbb{R}^{n}\times\mathbb{R}^{q}$ and $(\theta_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{q}$ such that $x^{*}\in R(\theta^{*})$ and $\theta_{k}\xrightarrow[k\to\infty]{}\theta^{*}$, there exists $(x_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{n}$ such that $x_{k}\xrightarrow[k\to\infty]{}x^{*}$ and $x_{k}\in R(\theta_{k})$ for any $k\in\mathbb{N}$.
Let $(x^{*},\theta^{*})\in\mathbb{R}^{n}\times\mathbb{R}^{q}$ such that $x^{*}\in R(\theta^{*})$, and $(\theta_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{q}$ such that $\theta_{k}\xrightarrow[k\to\infty]{}\theta^{*}$. Since $(x^{*},\theta^{*})$ is MFCQ-qualified, [32] proved that $R$ has the Aubin property, or equivalently, is pseudo-Lipschitzian, around $(x^{*},\theta^{*})$. That means there exist a neighborhood $\mathcal{N}$ of $x^{*}$, a neighborhood $\mathcal{U}$ of $\theta^{*}$ and $M\geq 0$ such that
@@
In particular, for any $k\geq K$, with $K$ large enough,
@@
Therefore, for any $k\geq K$ , there exists $x_{k}\in R(\theta_{k})$ and $z_{k}\in\mathcal{B}_{c}(0,1)$ such that
@@
The set-valued map of primal solutions $\mathrm{Sol_{p}}$ has a closed graph.
Let $(\theta_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{q}$, $(x_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{n}$ and $(x^{*},\theta^{*})\in\mathbb{R}^{n}\times\mathbb{R}^{q}$ such that $x_{k}\in\mathrm{Sol_{p}}(\theta_{k})$ for all $k\in\mathbb{N}$, $x_{k}\xrightarrow[k\to\infty]{}x^{*}$ and $\theta_{k}\xrightarrow[k\to\infty]{}\theta^{*}$. It’s immediate that $x^{*}\in R(\theta^{*})$.
@@
We can conclude that $x^{*}\in\mathrm{Sol_{p}}(\theta^{*})$ and that $\mathrm{Sol_{p}}$ is graph closed. ∎ The following lemma can also be obtained through the Berge’s maximum theorem.
@@
The value function $f$, defined in (1.1), is continuous on $\mathbb{R}^{q}$.
@@
For $K$ large enough, we have $\theta_{k}\in\mathcal{N}$ and $x_{k}\in\mathrm{Sol_{p}}(\theta_{k})\subset\mathcal{B}_{c}(0,r)$ whenever $k\geq K$. We then get that $(x_{k})_{k\in\mathbb{N}}$ is bounded. Therefore $(f(\theta_{k}))_{k\in\mathbb{N}}=(F(x_{k},\theta_{k}))_{k\in\mathbb{N}}$ is bounded by continuity of $F$.
@@
Lemma A.5 ([34, Lemma 9]).
Let Assumptions 4.2() hold. Let $D_{F}$ a be a definable conservative field for $F$. Then there exists a sequence of definable selections $\left\{(x_{\ell},\lambda_{\ell},\mu_{\ell})\right\}_{{\ell}\in\mathbb{N}}$ such that, for any $\ell$$\in\mathbb{N}$ and $\theta\in\mathbb{R}^{q}$, $(x_{\ell}(\theta),\lambda_{\ell}(\theta),\mu_{\ell}(\theta))\in\mathrm{Sol_{pd}}(\theta)$ and such that, for all $\theta\in\mathbb{R}^{q}$,
$\Big\{u\in\mathbb{R}^{q}:\exists\,(x,\lambda,\mu)\in\mathrm{Sol_{pd}}(\theta),(0,u)\in\operatorname*{conv\!}D_{F}(x,\theta)$
$\qquad\quad+\left(\operatorname{Jac}_{x}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{x}\!H(x,\theta)^{\!\mathsf{T}}\mu,\operatorname{Jac}_{\theta}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{\theta}\!H(x,\theta)^{\!\mathsf{T}}\mu\right)\Big\}$
@@
$\qquad\quad+\big(\operatorname{Jac}_{x}\!G(x_{\ell}(\theta),\theta)^{\!\mathsf{T}}\lambda_{\ell}(\theta)+\operatorname{Jac}_{x}\!H(x_{\ell}(\theta),\theta)^{\!\mathsf{T}}\mu_{\ell}(\theta),\operatorname{Jac}_{\theta}\!G(x_{\ell}(\theta),\theta)^{\!\mathsf{T}}\lambda_{\ell}(\theta)$
$\qquad\qquad\qquad+\operatorname{Jac}_{\theta}\!H(x_{\ell}(\theta),\theta)^{\!\mathsf{T}}\mu_{\ell}(\theta)\big)\big\}.$
@@
Let $\bar{z}\in A$. Then there exists $(c_{r_{s}})_{s\in\mathbb{N}}\subset A$ such $c_{r_{s}}\xrightarrow[s\to\infty]{}\bar{z}$. Let $\varepsilon>0$ and $N\in\mathbb{N}$. There exist $s\in\mathbb{N}$ such that $\lVert{{c_{r_{s}}}-{\bar{z}}}\rVert<\varepsilon/2$. Let $L\in\mathbb{N}$ such that $1/L<\varepsilon/2$ and $L>r_{s}$. We have $w_{\ell,r_{s}}\in\mathcal{B}(c_{r_{s}},1/\ell)$ for all $\ell\geq L$. It follows that $\lVert{{w_{\ell,r_{s}}}-{\bar{z}}}\rVert<\varepsilon$ for all $\ell\geq L$. Let $\ell\geq\max\{L,N\}$. By definition, there exists $k\geq N$ such that $z_{k}=w_{\ell,r_{s}}$. Then $\lVert{{z_{k}}-{\bar{z}}}\rVert<\varepsilon$ and $\bar{z}$ is an accumulation point of $(z_{k})_{k\in\mathbb{N}}$.
@@
This is a contradiction. So for every $\ell\geq L$, $\lVert{{\bar{z}}-{w_{\ell,s}}}\rVert\geq 1/L$ for all $s\in\{1,\ldots,\ell\}$. That means that there exists $K\in\mathbb{N}$ such that for any $k\geq K$, $\lVert{{z_{k}}-{\bar{z}}}\rVert\geq 1/L$. Thus $\bar{z}$ is not an accumulation point of $(z_{k})_{k\in\mathbb{N}}$.
@@
Let $C\subset\mathbb{R}^{\ell}$ be closed with empty interior. Then, there exists a differentiable and locally Lipschitz function $\bar{g}\colon\mathbb{R}^{\ell}\to\mathbb{R}$ such that $\bar{g}(z)=0$ for every $z\in C$, and $\bar{g}(z)>0$ for every $z\in\mathbb{R}^{\ell}\setminus C$. Moreover $\bar{g}$ is continuously differentiable everywhere except on $C$ with $\mathcal{B}_{c}(0,1)\subset\partial^{c}\bar{g}(z)$ for any $z\in C$. In particular, if $C$ is defined as in Section 4.4(), $\bar{g}$ is also path differentiable.
@@
By Whitney theorem [30, Theorem 2.29], $C$ is the zero of an infinitely differentiable and non-negative function $\bar{f}\colon\mathbb{R}^{\ell}\to\mathbb{R}$ and, for any $z\in C$, $\nabla\bar{f}(z)=0$. Thanks to Lemma A.6(), there exists a sequence $(z_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{2}\setminus C$ such that its set of accumulation points is $C$. For any $k\in\mathbb{N}$, consider $\mathcal{B}(z_{k},r_{k})$ such that $r_{k}>0$, $r_{k}\xrightarrow[k\to\infty]{}0$, and $\mathcal{B}(z_{k},r_{k})\cap\mathcal{B}(z_{s},r_{s})=\varnothing$ for $s\neq k$. Such a choice is possible, because for each $k$, $z_{k}$ is at positive distance of $z_{i}$ for each $i\neq k$ and at positive distance of $C$ the set of accumulation points. Define $\bar{g}\colon\mathbb{R}^{\ell}\to\mathbb{R}$ as
@@
Then, $\bar{g}(z)=0$ for every $z\in C$, and $\bar{g}(z)>0$ for every $z\in\mathbb{R}^{\ell}\setminus C$. Moreover $\bar{g}$ is differentiable and locally Lipschitz on $\mathbb{R}^{\ell}$. It is continuously differentiable everywhere except on $C$ because $\mathcal{B}_{c}(0,1)\subset\partial^{c}\bar{g}(z)$ for any $z\in C$. Indeed, let $\bar{z}\in C$ and, without loss of generality, we assume that $z_{k}\xrightarrow[k\to\infty]{}\bar{z}$. Let $v\in\{w\in\mathbb{R}^{\ell}:\lVert{w}\rVert=1\}$. Consider $(\bar{z}_{k})_{k\in\mathbb{N}}$ such that $\bar{z}_{k}=z_{k}-\frac{r_{k}}{2}v$ for all $k\in\mathbb{N}$. Then we have, for any $k\in\mathbb{N}$, $\nabla\bar{g}(\bar{z}_{k})=\nabla\bar{f}(\bar{z_{k}})+v$. So, $\bar{z}_{k}\xrightarrow[k\to\infty]{}\bar{z}$ and $\nabla\bar{g}(\bar{z}_{k})\xrightarrow[k\to\infty]{}v$. It follows that $v\in\partial^{c}\bar{g}(\bar{z})$. Since $v$ is arbitrary in $\{w\in\mathbb{R}^{\ell}:\lVert{w}\rVert=1\}$, and by Definition 3.1(), we have
$\operatorname*{conv\!}\,\{w\in\mathbb{R}^{\ell}:\lVert{w}\rVert=1\}=\mathcal{B}_{c}(0,1)\subset\partial^{c}\bar{g}(z).$
Now let us assume that $C$ is defined as in Section 4.4(). Let $\gamma\colon[0,1]\to\mathbb{R}^{\ell}$ be an absolutely continuous curve. Given that $\bar{g}$ is locally Lipschitz, $F\circ\gamma$ is absolutely continuous. Let $\Omega\subset[0,1]$ be the full measure set where $F\circ\gamma$ and $\gamma$ are differentiable. Consider the following three sets:
$S=\{t\in\Omega:\gamma(t)\in\mathbb{R}^{\ell}\setminus C\},$
$S_{1}=\{t\in\Omega:\gamma(t)\in C\text{ and }\dot{\gamma}(t)=0\},$
$S_{2}=\{t\in\Omega:\gamma(t)\in C\text{ and }\dot{\gamma}(t)\neq 0\}.$
We have $\Omega=S\cup S_{1}\cup S_{2}$. Since $\bar{g}$ is continuously differentiable outside $C$, the chain rule holds on $S$, i.e., for any $t\in S$,
@@
Thanks to [34, Lemma 6], we know that $S_{2}$ has zero measure. As $\bar{g}$ is locally Lipschitz and $\gamma(t+h)=\gamma(t)+o(h)$ for all $(t+h,t)\in\Omega\times S_{1}$ with $|h|$ small enough, the chain rule holds in $S_{1}$. In conclusion, the chain rule holds for almost all $t\in[0,1]$ and $\bar{g}$ is path differentiable. ∎
@@ Acknowledgements.
This work has been supported by the Occitanie region, the European Regional Development Fund (ERDF), and the French government, through the France 2030 project managed by the National Research Agency (ANR) with the reference number “ANR-22-EXES-0015.

### 22 Apr 2026 v2

@@ Appendix A Technical results
The following lemma is an immediate consequence of the stronger Aubin property result given in [38, Corollary 4.4]. We give its quick derivation in the proof just for completeness.
@@
Under Assumption 4.4(), the set-valued map of feasible points $R$ is inner semicontinuous, i.e., for all $(x^{*},\theta^{*})\in\mathbb{R}^{n}\times\mathbb{R}^{q}$ and $(\theta_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{q}$ such that $x^{*}\in R(\theta^{*})$ and $\theta_{k}\xrightarrow[k\to\infty]{}\theta^{*}$, there exists $(x_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{n}$ such that $x_{k}\xrightarrow[k\to\infty]{}x^{*}$ and $x_{k}\in R(\theta_{k})$ for any $k\in\mathbb{N}$.
Let $(x^{*},\theta^{*})\in\mathbb{R}^{n}\times\mathbb{R}^{q}$ such that $x^{*}\in R(\theta^{*})$, and $(\theta_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{q}$ such that $\theta_{k}\xrightarrow[k\to\infty]{}\theta^{*}$. Since $(x^{*},\theta^{*})$ is RCQ-qualified as in Definition 4.1(), [38, Corollary 4.4, (4.16)] ensures that $R$ has the Aubin property, or equivalently, is pseudo-Lipschitzian, around $(x^{*},\theta^{*})$. That means there exist a neighborhood $\mathcal{N}$ of $x^{*}$, a neighborhood $\mathcal{U}$ of $\theta^{*}$ and $M\geq 0$ such that
@@
In particular, for any $k\geq\bar{k}$, with $\bar{k}$ large enough,
@@
Therefore, for any $k\geq\bar{k}$ , there exists $x_{k}\in R(\theta_{k})$ and $z_{k}\in\mathcal{B}_{c}(0,1)$ such that
@@
Under Assumption 4.4(), the set-valued map of primal solutions $\mathrm{Sol_{p}}$ has a closed graph.
Let $(\theta_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{q}$, $(x_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{n}$ and $(x^{*},\theta^{*})\in\mathbb{R}^{n}\times\mathbb{R}^{q}$ such that $x_{k}\in\mathrm{Sol_{p}}(\theta_{k})$ for all $k\in\mathbb{N}$, $x_{k}\xrightarrow[k\to\infty]{}x^{*}$ and $\theta_{k}\xrightarrow[k\to\infty]{}\theta^{*}$. Using continuity of $C$ and closedness of $K$, we have that $x^{*}\in R(\theta^{*})$.
@@
We can conclude that $x^{*}\in\mathrm{Sol_{p}}(\theta^{*})$ and that $\mathrm{Sol_{p}}$ is graph closed. ∎ The following lemma can also be obtained through the Berge’s maximum theorem. We provide a self contained proof which fits Assumption 4.4()
@@
Under Assumption 4.4(), the value function $f$, defined in (4.1), is continuous on $\mathbb{R}^{q}$.
@@
For $\bar{k}$ large enough, we have $\theta_{k}\in\mathcal{N}$ and $x_{k}\in\mathrm{Sol_{p}}(\theta_{k})\subset\mathcal{B}_{c}(0,r)$ whenever $k\geq\bar{k}$. We then get that $(x_{k})_{k\in\mathbb{N}}$ is bounded. Therefore $(f(\theta_{k}))_{k\in\mathbb{N}}=(F(x_{k},\theta_{k}))_{k\in\mathbb{N}}$ is bounded by continuity of $F$.
@@
Lemma A.5 ([40, Lemma 9]).
Let Assumptions 4.4() hold. Let $D_{F}$ be a definable conservative field for $F$. Then there exists a sequence of definable selections $\left\{(x_{\ell},\lambda_{\ell})\right\}_{{\ell}\in\mathbb{N}}$ such that, for any $\ell$$\in\mathbb{N}$ and $\theta\in\mathbb{R}^{q}$, $(x_{\ell}(\theta),\lambda_{\ell}(\theta))\in\mathrm{Sol_{pd}}(\theta)$ and such that, for all $\theta\in\mathbb{R}^{q}$,
$\Big\{u\in\mathbb{R}^{q}:\exists\,(x,\lambda)\in\mathrm{Sol_{pd}}(\theta),(0,u)\in\operatorname*{conv\!}D_{F}(x,\theta)+\left(\operatorname{Jac}_{x}\!C(x,\theta)^{\!\mathsf{T}}\lambda,\operatorname{Jac}_{\theta}\!C(x,\theta)^{\!\mathsf{T}}\lambda\right)\Big\}$
@@
$\qquad\quad+\big(\operatorname{Jac}_{x}\!C(x_{\ell}(\theta),\theta)^{\!\mathsf{T}}\lambda_{\ell}(\theta),\operatorname{Jac}_{\theta}\!C(x_{\ell}(\theta),\theta)^{\!\mathsf{T}}\lambda_{\ell}(\theta)\big)\big\}.$
The following results are technical elements for the construction of the counter example of Section 5().
@@
Let $\bar{z}\in A$. Then there exists $(c_{r_{s}})_{s\in\mathbb{N}}\subset A$ such $c_{r_{s}}\xrightarrow[s\to\infty]{}\bar{z}$. Let $\varepsilon>0$ and $N\in\mathbb{N}$. There exists $s\in\mathbb{N}$ such that $\lVert{{c_{r_{s}}}-{\bar{z}}}\rVert<\varepsilon/2$. Let $L\in\mathbb{N}$ such that $1/L<\varepsilon/2$ and $L>r_{s}$. We have $w_{\ell,r_{s}}\in\mathcal{B}(c_{r_{s}},1/\ell)$ for all $\ell\geq L$. It follows that $\lVert{{w_{\ell,r_{s}}}-{\bar{z}}}\rVert<\varepsilon$ for all $\ell\geq L$. Let $\ell\geq\max\{L,N\}$. By definition, there exists $k\geq N$ such that $z_{k}=w_{\ell,r_{s}}$. Then $\lVert{{z_{k}}-{\bar{z}}}\rVert<\varepsilon$ and $\bar{z}$ is an accumulation point of $(z_{k})_{k\in\mathbb{N}}$.
@@
This is a contradiction. So for every $\ell\geq L$, $\lVert{{\bar{z}}-{w_{\ell,s}}}\rVert\geq 1/L$ for all $s\in\{1,\ldots,\ell\}$. That means that there exists $\bar{k}\in\mathbb{N}$ such that for any $k\geq\bar{k}$, $\lVert{{z_{k}}-{\bar{z}}}\rVert\geq 1/L$. Thus $\bar{z}$ is not an accumulation point of $(z_{k})_{k\in\mathbb{N}}$.
@@
Let $S\subset\mathbb{R}^{\ell}$ be closed with empty interior. Then, there exists a differentiable and locally Lipschitz function $\bar{g}\colon\mathbb{R}^{\ell}\to\mathbb{R}$ such that $\bar{g}(z)=0$ for every $z\in S$, and $\bar{g}(z)>0$ for every $z\in\mathbb{R}^{\ell}\setminus S$. Moreover $\bar{g}$ is continuously differentiable everywhere except on $S$ with $\mathcal{B}_{c}(0,1)\subset\partial^{c}\bar{g}(z)$ for any $z\in S$. In particular, if $S$ is defined as in Section 5(), $\bar{g}$ is also path differentiable.
@@
By Whitney theorem [36, Theorem 2.29], $S$ is the zero of an infinitely differentiable and non-negative function $\bar{f}\colon\mathbb{R}^{\ell}\to\mathbb{R}$ and, for any $z\in S$, $\nabla\bar{f}(z)=0$. Thanks to Lemma A.6(), there exists a sequence $(z_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{\ell}\setminus S$ such that its set of accumulation points is $S$. For any $k\in\mathbb{N}$, consider $\mathcal{B}(z_{k},r_{k})$ such that $r_{k}>0$, $r_{k}\xrightarrow[k\to\infty]{}0$, and $\mathcal{B}(z_{k},r_{k})\cap\mathcal{B}(z_{s},r_{s})=\varnothing$ for $s\neq k$. Such a choice is possible, because for each $k$, $z_{k}$ is at positive distance of $z_{i}$ for each $i\neq k$ and at positive distance of $S$ the set of accumulation points. Define $\bar{g}\colon\mathbb{R}^{\ell}\to\mathbb{R}$ as
@@
Then, $\bar{g}(z)=0$ for every $z\in S$, and $\bar{g}(z)>0$ for every $z\in\mathbb{R}^{\ell}\setminus S$. Moreover $\bar{g}$ is differentiable and locally Lipschitz on $\mathbb{R}^{\ell}$. It is continuously differentiable everywhere except on $S$ because $\mathcal{B}_{c}(0,1)\subset\partial^{c}\bar{g}(z)$ for any $z\in S$. Indeed, let $\bar{z}\in S$ and, without loss of generality, we assume that $z_{k}\xrightarrow[k\to\infty]{}\bar{z}$. Let $v\in\{w\in\mathbb{R}^{\ell}:\lVert{w}\rVert=1\}$. Consider $(\bar{z}_{k})_{k\in\mathbb{N}}$ such that $\bar{z}_{k}=z_{k}-\frac{r_{k}}{2}v$ for all $k\in\mathbb{N}$. Then we have, for any $k\in\mathbb{N}$, $\nabla\bar{g}(\bar{z}_{k})=\nabla\bar{f}(\bar{z_{k}})+v$. So, $\bar{z}_{k}\xrightarrow[k\to\infty]{}\bar{z}$ and $\nabla\bar{g}(\bar{z}_{k})\xrightarrow[k\to\infty]{}v$. It follows that $v\in\partial^{c}\bar{g}(\bar{z})$. Since $v$ is arbitrary in $\{w\in\mathbb{R}^{\ell}:\lVert{w}\rVert=1\}$, and by Definition 3.1(), we have
$\operatorname*{conv\!}\,\{w\in\mathbb{R}^{\ell}:\lVert{w}\rVert=1\}=\mathcal{B}_{c}(0,1)\subset\partial^{c}\bar{g}(\bar{z}).$
Now let us assume that $S$ is defined as in Section 5(). Let $\gamma\colon[0,1]\to\mathbb{R}^{\ell}$ be an absolutely continuous curve. Given that $\bar{g}$ is locally Lipschitz, $\bar{g}\circ\gamma$ is absolutely continuous. Let $\Omega\subset[0,1]$ be the full measure set where $\bar{g}\circ\gamma$ and $\gamma$ are differentiable. Consider the following three sets:
$S_{1}=\{t\in\Omega:\gamma(t)\in\mathbb{R}^{\ell}\setminus S\},$
$S_{2}=\{t\in\Omega:\gamma(t)\in S\text{ and }\dot{\gamma}(t)=0\},$
$S_{3}=\{t\in\Omega:\gamma(t)\in S\text{ and }\dot{\gamma}(t)\neq 0\}.$
We have $\Omega=S_{1}\cup S_{2}\cup S_{3}$. Since $\bar{g}$ is continuously differentiable outside $S$, the chain rule holds on $S_{1}$, i.e., for any $t\in S_{1}$,
@@
Thanks to [40, Lemma 6], we know that $S_{3}$ has zero measure. As $\bar{g}$ is locally Lipschitz and $\gamma(t+h)=\gamma(t)+o(h)$ for all $(t+h,t)\in\Omega\times S_{2}$ with $|h|$ small enough, the chain rule holds in $S_{2}$. In conclusion, the chain rule holds for almost all $t\in[0,1]$ and $\bar{g}$ is path differentiable. ∎
@@ Acknowledgements.
This work has been supported by the Occitanie region, the European Regional Development Fund (ERDF), and the French government, through the France 2030 project managed by the National Research Agency (ANR) with the reference number “ANR-22-EXES-0015”.